# Into the vortex: quasinormal and quasibound spectra of rotating acoustic geometries

### v1

@@ I Introduction
Analog black-hole (BH) experiments [1, 2] are currently playing an important role towards a more complete understanding of classical and quantum effects around astrophysical BHs [3]. From such gravity simulators, we have so far obtained information regarding BH spectroscopy [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], superradiance [30, 31, 32, 33, 34, 35, 36, 37], Hawking radiation [38, 39, 40, 41, 42, 43, 44], and BH evaporation [45, 46, 47]. These research topics have either recently begun to emerge through gravitational-wave (GW) detections [48, 49] or have not yet been explored in fully relativistic astrophysical setups [50, 51, 52, 53, 54]. Ultimately, we expect that the critical prognosis provided by various analog gravity simulations, will further elucidate the classical, semi-classical [55, 56, 57] and quantum nature of spacetime [58, 59].
@@
Currently, a multitude of gravity simulators have been built using various fluids [60, 61, 62, 41, 34, 63, 33, 14, 64, 13, 10, 65, 66, 67, 68], Bose-Einstein condensates [69, 70, 71, 72, 73, 74, 75, 76, 77], lasers [78, 79, 80, 43, 81, 82], optical media [83, 84, 35, 85, 86], and electric circuits [87, 88], to name a few. Even though the above experiments have reproduced both realistic and theoretical phenomenology of BHs and compact objects, all experimental setups can experience instrumental limitations (water tank size, fluid dissipation [68], bubble-forming instability at high rotation rates [89], thermal equilibrium of semiconductor lasers [90], only classical BH effects for optical experiments [86], negative-refraction metamaterials [88, 91], dispersion [92]). Thus, the more experiments performed with different instrumentation, the better we will understand the effectiveness of BH analogs. For example, the idea of using different states of matter, such as superconductivity [68, 93, 94] and superfluidity [68, 92], and even ferromagnetism [95], enables a more in-depth perception of analog systems, due to their efficient experimental properties. One very effective example of a state of matter that directly connects to experiments is superfluidity. This is a characteristic state of matter of a fluid when it undergoes a phase transition at low-temperatures and leads to a material that has zero viscosity and, therefore, flows without any loss of kinetic energy. Thus, when stirred, a vortex forms that rotate indefinitely which can be used to study quantum effects related to event horizons, superradiance and ergoregions [68, 96, 97, 98, 99], and even the dynamics of the early Universe [100, 101].
@@ III.1 Quasinormal modes
The behavior of corotating and counterrotating QNMs is illustrated in Fig. 1 as a function of the tuning parameter $a$ and the circulation $C$. A universal trend is found for various values of $C$, where both the real and the absolute value of the imaginary part of QNMs increase with the tuning parameter $a$. A symmetry around $C=0$ is observed for corotating and counterrotating modes. This is merely due to the role that $C$ plays in the potential (26). When $C$ is set to zero, all terms that are multiplied with it, such as $m$, play no role in the dynamics of the potential, no matter the magnetic number. The QNMs provided here can, in principle, shed more light in experiments such as the one in [92], where evidence of BH ringdown counterrotating modes ($m<0$) have been observ in hollow-core superfluid 4He vortices with circulation.
@@ Appendix A Quasibound states with the VBK approach
By substituting Eqs. (34)-(38) into Eq. (LABEL:eq:radial_equation_UABHc_2), we get
@@
The radial equation (39) has the form of a confluent Heun equation [149], which is given by

### v2

@@ I Introduction
Analog black-hole (BH) experiments [1, 2, 3] are currently playing an important role towards a more complete understanding of classical and quantum effects around astrophysical BHs [4]. From such gravity simulators, we have so far obtained information regarding BH spectroscopy [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30], superradiance [31, 32, 33, 34, 35, 36, 37, 38], Hawking radiation [39, 40, 41, 42, 43, 44, 45], and BH evaporation [46, 47, 48]. These research topics have either recently begun to emerge through gravitational-wave (GW) detections [49, 50] or have not yet been explored in fully relativistic astrophysical setups [51, 52, 53, 54, 55]. Ultimately, we expect that the critical prognosis provided by various analog gravity simulations, will further elucidate the classical, semi-classical [56, 57, 58] and quantum nature of spacetime [59, 60].
@@
Currently, a multitude of gravity simulators have been built using various fluids [61, 62, 63, 42, 35, 64, 34, 15, 65, 14, 11, 66, 67, 68, 69], Bose-Einstein condensates [70, 71, 72, 73, 74, 75, 76, 77, 78], lasers [79, 80, 81, 44, 82, 83], optical media [84, 85, 36, 86, 87], and electric circuits [88, 89], to name a few. Even though the above experiments have reproduced both realistic and theoretical phenomenology of BHs and compact objects, all experimental setups can experience instrumental limitations (water tank size, fluid dissipation [69], bubble-forming instability at high rotation rates [90], thermal equilibrium of semiconductor lasers [91], only classical BH effects for optical experiments [87], negative-refraction metamaterials [89, 92], dispersion [93]). Thus, the more experiments performed with different instrumentation, the better we will understand the effectiveness of BH analogs. For example, the idea of using different states of matter, such as superconductivity [69, 94, 95] and superfluidity [69, 93], Bose-Einstein condensates [96, 97, 98], and even ferromagnetism [99], enables a more in-depth perception of analog systems, due to their efficient experimental properties. One very effective example of a state of matter that directly connects to experiments is superfluidity. This is a characteristic state of matter of a fluid when it undergoes a phase transition at low-temperatures and leads to a material that has zero viscosity and, therefore, flows without any loss of kinetic energy. Thus, when stirred, a vortex forms that rotate indefinitely which can be used to study quantum effects related to event horizons, superradiance and ergoregions [69, 100, 101, 102, 103], and even the dynamics of the early Universe and cosmology [104, 105, 106, 107, 108].
@@ III.1 Quasinormal modes
The behavior of corotating and counterrotating QNMs is illustrated in Fig. 1 as a function of the tuning parameter $a$ and the circulation $C$. A universal trend is found for various values of $C$, where both the real and the absolute value of the imaginary part of QNMs increase with the tuning parameter $a$. A symmetry around $C=0$ is observed for corotating and counterrotating modes. This is merely due to the role that $C$ plays in the potential (26). When $C$ is set to zero, all terms that are multiplied with it, such as $m$, play no role in the dynamics of the potential, no matter the magnetic number. The QNMs provided here can, in principle, shed more light in experiments such as the one in [93], where evidence of BH ringdown counterrotating modes ($m<0$) have been observed in hollow-core superfluid 4He vortices with circulation.
@@ Appendix A Quasibound states with the VBK approach
By substituting Eqs. (34)-(38) into Eq. (24), we get
@@
The radial equation (39) has the form of a confluent Heun equation [156, 157], which is given by